$A$ die is thrown $1000$ times with the frequencies for the outcomes $1, 2, 3, 4, 5$ and $6$ as given in the following table:
Outcome $1$ $2$ $3$ $4$ $5$ $6$
Frequency $179$ $150$ $157$ $149$ $175$ $190$

Find the probability of getting each outcome.

Vedclass pdf generator app on play store
Vedclass iOS app on app store
(N/A) Let $E_i$ denote the event of getting the outcome $i$,where $i = 1, 2, 3, 4, 5, 6$.
The probability of an event is given by the formula: $P(E) = \frac{\text{Number of trials in which the event happened}}{\text{Total number of trials}}$.
Here,the total number of trials is $1000$.
$P(E_1) = \frac{179}{1000} = 0.179$
$P(E_2) = \frac{150}{1000} = 0.150$
$P(E_3) = \frac{157}{1000} = 0.157$
$P(E_4) = \frac{149}{1000} = 0.149$
$P(E_5) = \frac{175}{1000} = 0.175$
$P(E_6) = \frac{190}{1000} = 0.190$
Note that the sum of all probabilities is $0.179 + 0.150 + 0.157 + 0.149 + 0.175 + 0.190 = 1$.

Explore More

Similar Questions

Find the probability that a student of the class was born in August.

$A$ coin is tossed $1000$ times with the following frequencies:
Head : $455$ Tail : $545$
Compute the probability for each event.

The distance (in $km$) of $40$ engineers from their residence to their place of work were found as follows.
$5$ $3$ $10$ $20$ $25$ $11$ $13$ $7$ $12$ $31$
$19$ $10$ $12$ $17$ $18$ $11$ $32$ $17$ $16$ $2$
$7$ $9$ $7$ $8$ $3$ $5$ $12$ $15$ $18$ $3$
$12$ $14$ $2$ $9$ $6$ $15$ $15$ $7$ $6$ $12$

What is the empirical probability that an engineer lives:
$(i)$ less than $7 \, km$ from her place of work?
$(ii)$ more than or equal to $7 \, km$ from her place of work?
$(iii)$ within $\frac{1}{2} \, km$ from her place of work?

The following frequency distribution table shows the concentration of sulphur dioxide $(SO_2)$ in the air in parts per million $(ppm)$ of a certain city for $30$ days. Using this table,find the probability of the concentration of sulphur dioxide being in the interval $0.12 - 0.16$ on any of these days.
Concentration of $SO_2$ (in $ppm$)Number of days (frequency)
$0.00-0.04$$4$
$0.04-0.08$$9$
$0.08-0.12$$9$
$0.12-0.16$$2$
$0.16-0.20$$4$
$0.20-0.24$$2$
Total$30$

$A$ teacher wanted to analyze the performance of two sections of students in a mathematics test of $100$ marks. Looking at their performances,she found that a few students got under $20$ marks and a few got $70$ marks or above. So she decided to group them into intervals of varying sizes as follows: $0-20, 20-30, ..., 60-70, 70-100$. Then she formed the following table:
Marks Number of students
$0-20$ $7$
$20-30$ $10$
$30-40$ $10$
$40-50$ $20$
$50-60$ $20$
$60-70$ $15$
$70$ and above $8$
Total $90$

$(i)$ Find the probability that a student obtained less than $20\%$ in the mathematics test.
$(ii)$ Find the probability that a student obtained marks $60$ or above.

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo